OpenStudy (anonymous):

Divide. 4/5÷2/3 A.8/15 B.5/6 C.1 1/5 D.1 7/8

OpenStudy (anonymous):

\[\huge \frac{4}{5}\div \frac{2}{3}=\frac{4}{5}\times \frac{3}{2}\] cancel first, multiply last

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

is it clear how to multiply those two fractions?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ok you should get \(\frac{6}{5}\) when you are done then turn in to a mixed number

OpenStudy (anonymous):

Its my little sisters math problem.

OpenStudy (anonymous):

lol tell her "when in doubt, Charlie out"

OpenStudy (just_one_last_goodbye):

Well I wanted to help but... everyone is attacking this question as if it was the last $1 Ipad for sale on black friday O_O

OpenStudy (compassionate):

@satellite73 , I think you mean, "Cross multiply." Not multiply.

OpenStudy (just_one_last_goodbye):

^ actually no its not that Comp

OpenStudy (just_one_last_goodbye):

\[\frac{ 3 }{ 2 } * \frac{ 4 }{ 5 } = \frac{ 12 }{ 10 }\]

OpenStudy (anonymous):

dude its 1 1/5

OpenStudy (just_one_last_goodbye):

^ now simplify \[\frac{ 12 }{ 10 }\]

OpenStudy (anonymous):

yes it is \(1\tfrac{1}{5}\)

OpenStudy (anonymous):

@just_one_last_goodbye there is no such mathematical operation as "simplify" @Compassionate there is no such mathematical operation as "cross multiply"

OpenStudy (compassionate):

Oh, so @e.mccormick was right. You really do hate the term "cross multiply." Thank you for proving that theory.

OpenStudy (anonymous):

i don't hate it like i don't hate a unicorn it doesn't exist

OpenStudy (compassionate):

Cross multiplication confirmed for unicornary